OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 − 2x + 4 8x + y = 20 (2, 84) and (−8, 4) (−8, 2) and (84, 4) (2, 4) and (−8, 84) (4, −8) and (84, 2)

OpenStudy (anonymous):

No, Mac. @mcfizzy321

rishavraj (rishavraj):

u can write the 2nd equation as y = 20 - 8x

OpenStudy (usukidoll):

mcfizzy, stop with these programs. Uploading malicious content will only get you into further trouble just like yesterday . geez.

rishavraj (rishavraj):

as y = x^2 - 2x + 4 and y = 20 - 8x so x^2 - 2x + 4 = 20 - 8x proceed

OpenStudy (anonymous):

So, would combine x2 and 2x and then continue with simplifying?

rishavraj (rishavraj):

\[x^2 - 2x + 4 = 20 - 8x\] \[x^2 + 6x - 16 = 0\] solve the equation

OpenStudy (anonymous):

Okay, so the factors would be (x-8)(x+2)... is this right so far?

rishavraj (rishavraj):

it would be (x + 8)(x - 2)

OpenStudy (anonymous):

Okay, got it. Is this it, or is there more to do? Sorry, I was so confused on this one.

OpenStudy (usukidoll):

I'm wondering... how do we find (84,4) ?

OpenStudy (anonymous):

I have no idea, that's what I was confused about too.

OpenStudy (usukidoll):

because (x+8)(x-2) we have x = -8, 2 ... plug them in and solve for y?!

OpenStudy (usukidoll):

wait I think so... like if x = 2 8(2) + y = 20 16+y=20 y=20-16 = 4

OpenStudy (usukidoll):

duh I was getting there k...

OpenStudy (anonymous):

Ahhhh lol. Makes sense!

OpenStudy (anonymous):

So, it would be B?

OpenStudy (usukidoll):

we need our answer in (x,y) form like when x = 2, y = 4 so (2,4) was one pair x = -8, y = 84 (-8,84) was the other one

OpenStudy (anonymous):

Okay, so C?

OpenStudy (usukidoll):

yes

OpenStudy (anonymous):

Thank you so much!!!!